# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

### v2

@@ 4.2 Reasons for the failure of TS foundation models on DSR
Why do TS foundation models perform so poorly on DSR problems? Essentially, they face three key issues: First, most TS foundation models cannot efficiently deal with multivariate time series, but treat dimensions independently. However, in a nonlinear DS all variables are usually coupled, and their joint evolution is actually defining for the system’s (long-term) dynamics and attractor states. Thus, ignoring this, as illustrated in Fig. 6a, results in poor long-term accuracy (see also Figs. 16-21). This highlights a fundamental limitation: Univariate approaches are inherently insufficient for reconstructing multivariate, interacting DS. Second, TS foundation models are trained for short-term prediction, but not for DSR, providing another reason why they fail to reconstruct the correct long-term behavior. As illustrated in Figs. 1 & 6b, in the longer run they often converge to fixed points or simple cycles where the true dynamics is chaotic, not reflecting the context anymore. Third, context parroting, a phenomenon originally identified in LLMs [4], has also been observed in TS foundation models [99], where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution. Thus, TS foundation models simply produce oscillations, where the true dynamics is chaotic and therefore by definition irregular and non-repeating (Fig. 6c).
@@
This is illustrated in more detail in Fig. 6d for Chronos-t5-base, the best of the TS foundation models in our batch. It is particularly evident in the power spectrum (Fig. 6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix. We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [58], a measure for the exponential divergence rate of trajectories (see Appx. A.1). For Chronos, the value $\lambda_{Chronos}\approx 0.02$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 0.96$ close to the ground truth value of $\lambda_{Lorenz63}\approx 0.87$. Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context. On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig. 6e.
@@ 4.4 Time series forecasting from a DS perspective
We next tested and compared DynaMix on a variety of real-world datasets often used to probe time series models. As discussed in Sect. 3.3, for DynaMix we used an embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties (positional embedding for all empirical time series, akin to what is done in TS foundation models, and delay embedding for the Lorenz-63). Fig. 8a confirms for a $1d$ time series from the $3d$ chaotic Lorenz-63 system that this works well and DynaMix, in contrast to all other models, is still able to reconstruct the chaotic behavior. When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results [84], our model produces accurate forecasts which reflect the dynamics well (Fig. 8b-f, see Appx. A.2 for further details on the datasets). Our model seems to excel particularly on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig. 8g,h).
Table 1 quantitatively confirms that DynaMix mostly outperforms the other models tested on the DSR measures. Surprisingly, it often even outperforms them in terms of the forecasting error MAE, although its training corpus consisted purely of the simulated DS shown in Fig. 10, i.e. did not include any empirical data at all, let alone traffic, cloud usage, electricity, or temperature data! This is in stark contrast to the other TS foundation models tested here, which mostly had examples of such empirical data in their training corpus and hence might be expected to have an edge over DynaMix. Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos [58, 50], potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
@@ 4.5 Which ingredients are most crucial for DynaMix’ performance? Ablation studies
Which are the components of the DynaMix framework contributing most to its success in zero-shot out-of-domain generalization? Training by STF (Fig. 26), a sufficiently diverse training corpus, a sufficient number of experts (Fig. 25), the attention mechanism, and context preprocessing by the CNN, all appear vital to its strong performance (Fig. 9). Replacing the final gating step (the MLP) by a simple linear operation and the exact expert model within the class of piecewise linear RNNs, however, had less of an impact. Similar results were obtained when the AL-RNN was swapped for the clipped shallow PLRNN [39], and a probabilistic version of the AL-RNN (adding a noise term $\bm{z}_{t}=F(\bm{z}_{t-1})+\bm{\eta},\bm{\eta}\sim\mathcal{N}(0,\Gamma)$) even slightly improved long-term performance, albeit at the cost of some short-term prediction accuracy. Using as experts LSTMs [41], vanilla RNNs or reservoir computers [65] instead led to decay in performance, in line with previous results establishing PLRNNs as SOTA DSR models which allow for efficient training by STF [6, 39, 7]. Hence, it appears that the whole ‘DSR package’ consisting of 1) architectural choices, 2) a training algorithm specialized for DSR, and 3) properties of the DS training corpus, is crucial to DynaMix’ success.
@@ 5 Conclusions
Here we introduce the first DSR foundation model that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning. It hugely outperforms current TS foundation models like Chronos in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters. Consequently, DynaMix features inference times about an order of magnitude faster than the closest competitor. Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to its competitors – never seen in training, often producing also better short-term forecasts. All of this is achieved with a fairly narrow training corpus consisting mainly of $3d$ synthetic data from chaotic systems, which suggests there may be considerable room for further improvement by extending DynaMix’ training to include, e.g., various empirical time series.
One important take-home from this work therefore is that foundation models built based on principles of DS theory may be able to profoundly improve performance of current zero-shot TS forecasters. Most (if not all) empirically observed time series come from some underlying DS, and acknowledging this fact in model training and construction may help to advance the field. Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role. Most complex DS in nature [15, 86, 59], engineering [78], and society [56, 92] are likely chaotic. Moreover, many chaotic attractors, like the Lorenz-63, feature a skeleton of infinitely many (unstable) periodic orbits of all possible periods [36], thus inherently expressing a wide spectrum of temporal patterns.
@@
For now, we have mainly focused on rather stationary synthetic and empirical time series (but see Fig. 8b). Changes in statistical properties over time, tipping points, or widely differing time scales in the data, impose severe difficulties for zero-shot forecasting, as illustrated in Fig. 31a,b, which also plague other TS foundation models. This may partly be amended by explicitly including such types of non-stationary and multiscale DS in the training corpus. Another potential solution may be adding explicit filtering and decomposition modules, similar as in Auto- [95] or FEDformers [101]. A proof of concept of this idea is provided in Appx. A.5. Incorrect embeddings, which may not be immediately obvious, can also lead to forecast failures (Fig. 31c). Finally, continuous-time versions of the experts which can deal with sampling at irregular time points would be another fruitful direction.
@@ A.1 Methodological details
For training our model, we used a variant of sparse teacher forcing (STF), a control-theoretic method for DSR, with the theoretical background and rationale developed in [58]. STF and related training methods [6, 8, 39, 58, 24] circumvent the exploding gradient problem when training on chaotic systems, while still enabling the DSR model to sufficiently forward-iterate trajectories into the future to capture long-term properties (see [58] for details), and produces state-of-the-art results for DSR [7, 9]. In STF, current latent states states $\bm{z}_{t}$ are replaced with states inferred from the data at fixed intervals $\tau$ by (pseudo-)inversion of the decoder model, $\tilde{\bm{z}}_{t}=g^{-1}(\bm{x}_{t})$, thus recalibrating the trajectory at times ideally chosen based on the system’s Lyapunov spectrum [58] (or simply treating $\tau$ as a hyper-parameter). In our case of an identity mapping from a latent subspace to the observations, $\hat{\bm{x}}_{t}=\mathcal{I}_{(N\times M)}\bm{z}_{t}$, this inversion becomes trivial, where $N$ is the number of observations (and thus readout neurons) and $\mathcal{I}_{(N\times M)}$ an $N\times M$ matrix with $\mathcal{I}_{(N\times M),rr}=1$ for $r\leq N$ and zeros else. Thus, during training, we force
@@
where we chose $\lambda=0.1$ and $c=0.01$. Rectified adaptive moment estimation (RADAM) [52] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total. We used a learning rate exponentially decaying from $\eta_{\text{start}}=5\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$. The context length used in training was set to $T_{C}=500$, and the window of overlap with the model-generated time series to $\Delta t=50$, see Sect. 3.2. Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $30$ seconds depending on sequence length and model size. At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step. Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.1,0.2\}$, $\Delta t=\{0,50,80\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
@@
We use $J=10$ AL-RNN experts for our model. Each expert has a latent dimension of $M=30$, of which $P=2$ are rectified-linear units (ReLUs). As suggested in [7], we only use the first $N$ of the MPM-P linear units for the readout. We followed the initialization protocol in [6, 81], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix. The initial latent state was estimated as
@@
The gating network is implemented using a single-layer CNN with three channels and a kernel size of 2, stride of 1, and zero padding, with the identity as activation function. The MLP, which takes as inputs the concatenated weighted CNN outputs and current latent state, consists of two layers with ReLU activation. We initialize the temperature weights $\tau_{\text{att}}$ and $\tau_{\text{exp}}$ to $0.1$, the covariance matrix $\bm{\Sigma}=0.05\cdot\mathds{1}$, matrix $\bm{D}=\mathcal{I}_{(N\times M)}$, and draw all other CNN and MLP matrices from a Gaussian with standard deviation $0.01$. Model hyperparameters were selected according to previous results in [7] and by probing a few relevant parameter settings, namely $J=\{1,3,5,10,20,30\}$ (Fig. 25), $\text{CNN-layers}=\{1,2\}$, $\text{CNN-kernel-size}=\{2,3,5\}$, $\text{MLP-layers}=\{1,2,3\}$, and different activation functions $\{\mathds{1},\tanh{},\text{ReLU}\}$ for the MLP/CNN. We found that an extensive grid search was not necessary to obtain a well-performing model.
@@
As in [48, 58, 39, 8, 64, 30, 99], we assess the geometric (dis)similarity between true and model-generated attractors using a Kullback-Leibler divergence ($D_{\text{stsp}}$) defined across state space. Specifically, $D_{\text{stsp}}$ determines the overlap between the distributions of true trajectory points, $p_{true}(\bm{x})$, and of model-generated trajectories, $p_{gen}(\bm{x}|\bm{z})$, given by
@@
While especially for higher-dimensional state spaces these distributions may be approximated by Gaussian mixture models [48], here a discretized version was sufficient, with the state space parcellated into $K=m^{N}$ bins, where $m$ is the number of bins per dimension and $N$ the dimensionality of the system. We estimate the occupation probabilities $p_{i}$ of each bin via the relative frequencies $\hat{p}_{i}$ of trajectory visits and approximate $D_{\text{stsp}}$ as
@@
We set $m=30$ bins per dimension, following [38]. For the $6d$ Lorenz-96 we use $m=5$ and for the empirical time series $m=20$. To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled from the DS.
To evaluate the agreement in long-term temporal dynamics, we compute the Hellinger distance $D_{H}$ between the power spectra of the true and model-generated time series [58, 39], defined as
@@
where $F(\omega)$ and $G(\omega)$ are the power spectra of the true and generated time series, respectively. Power spectra are obtained via the dimension-wise Fast Fourier Transform (FFT), smoothed using a Gaussian kernel (with $\sigma=20$ for DS and $\sigma=2$ for the short empirical time series), and normalized to ensure comparability. High-frequency tails dominated by noise are truncated as described in [39]. The aggregated Hellinger distance $D_{H}$ is then computed as the average across all dimension-wise spectral comparisons.
For assessing short-term forecast quality aggregated across the whole test set of $54$ DS, we used a normalized $n$-step ahead prediction error (with $n=10$ in our evaluation) as recommended in [40], given by
@@
For individual comparisons on single empirical times, the MAE for evaluating short term predictions [40] was used:
@@
For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm [72]. First, the time series is embedded into a state space using time-delay embedding as in eq. 5 (e.g. using the PECUZAL algorithm [49]). For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [72]) and $l_{s}=0.25$. The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
@@
In chaotic systems, for a suitable choice of $k_{max}$ (here: $50$), this quantity initially grows approximately linearly with time [44], with the slope given by $\lambda_{\max}>0$. In contrast, periodic (cyclic) systems have $\lambda_{\max}=0$.

### v3

@@ 4.2 Reasons for the failure of TS foundation models on DSR
Why do TS foundation models perform so poorly on DSR problems? Essentially, they face three key issues: First, most TS foundation models cannot efficiently deal with multivariate time series, but treat dimensions independently. However, in a nonlinear DS all variables are usually coupled, and their joint evolution is actually defining for the system’s (long-term) dynamics and attractor states. Thus, ignoring this, as illustrated in Fig. 6a, results in poor long-term accuracy (see also Figs. 16-21). This highlights a fundamental limitation: Univariate approaches are inherently insufficient for reconstructing multivariate, interacting DS. Second, TS foundation models are trained for short-term prediction, but not for DSR, providing another reason why they fail to reconstruct the correct long-term behavior. As illustrated in Figs. 1 & 6b, in the longer run they often converge to fixed points or simple cycles where the true dynamics is chaotic, not reflecting the context anymore. Third, context parroting, a phenomenon originally identified in LLMs Bender et al. (2021), has also been observed in TS foundation models Zhang and Gilpin (2025), where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution. Thus, TS foundation models simply produce oscillations, where the true dynamics is chaotic and therefore by definition irregular and non-repeating (Fig. 6c).
@@
This is illustrated in more detail in Fig. 6d for Chronos-t5-base, the best of the TS foundation models according to $D_{stsp}$. It is particularly evident in the power spectrum (Fig. 6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix. We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent Mikhaeil et al. (2022), a measure for the exponential divergence rate of trajectories (see Appx. A.1). For Chronos, the value $\lambda_{Chronos}\approx 0.02$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 0.96$ close to the ground truth value of $\lambda_{Lorenz63}\approx 0.87$. Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context. On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig. 6e.
@@ 4.4 Time series forecasting from a DS perspective
We next tested and compared DynaMix on a variety of real-world datasets often used to probe time series models. As discussed in Sect. 3.3, for DynaMix we used an embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties (positional embedding for all empirical time series, akin to what is done in TS foundation models, and delay embedding for the Lorenz-63). Fig. 8a confirms for a $1d$ time series from the $3d$ chaotic Lorenz-63 system that this works well and DynaMix, in contrast to all other models, is still able to reconstruct the chaotic behavior. When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results Toner et al. (2025), our model produces accurate forecasts which reflect the dynamics well (Fig. 8b-f, see Appx. A.2 for further details on the datasets). Our model seems to excel particularly on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig. 8g,h).
Table 1 quantitatively confirms that DynaMix mostly outperforms the other models tested on the DSR measures. Surprisingly, it often even outperforms them in terms of the forecasting error MAE, although its training corpus consisted purely of the simulated DS shown in Fig. 10, i.e. did not include any empirical data at all, let alone traffic, cloud usage, electricity, or temperature data! This is in stark contrast to the other TS foundation models tested here, which mostly had examples of such empirical data in their training corpus and hence might be expected to have an edge over DynaMix. Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos Mikhaeil et al. (2022); Kramer et al. (2022), potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
@@ 4.5 Which ingredients are most crucial for DynaMix’ performance? Ablation studies
Which are the components of the DynaMix framework contributing most to its success in zero-shot out-of-domain generalization? Training by STF (Fig. 26), a sufficiently diverse training corpus, a sufficient number of experts (Fig. 25), the attention mechanism, and context preprocessing by the CNN, all appear vital to its strong performance (Fig. 9). Replacing the final gating step (the MLP) by a simple linear operation and the exact expert model within the class of piecewise linear RNNs, however, had less of an impact. Similar results were obtained when the AL-RNN was swapped for the clipped shallow PLRNN Hess et al. (2023), and a probabilistic version of the AL-RNN (adding a noise term $\bm{z}_{t}=F(\bm{z}_{t-1})+\bm{\eta},\bm{\eta}\sim\mathcal{N}(0,\Gamma)$) even slightly improved long-term performance, albeit at the cost of some short-term prediction accuracy. Using as experts LSTMs Hochreiter and Schmidhuber (1997), vanilla RNNs or reservoir computers Patel and Ott (2022) instead led to decay in performance, in line with previous results establishing PLRNNs as SOTA DSR models which allow for efficient training by STF Brenner et al. (2022); Hess et al. (2023); Brenner et al. (2024a). Hence, it appears that the whole ‘DSR package’ consisting of 1) architectural choices, 2) a training algorithm specialized for DSR, and 3) properties of the DS training corpus, is crucial to DynaMix’ success.
@@ 5 Conclusions
Here we introduce the first DSR foundation model that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning. It hugely outperforms current TS foundation models like Chronos or Panda in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters. Consequently, DynaMix features inference times about an order of magnitude faster than the closest competitor. Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to competitors – never seen in training, often producing also better short-term forecasts. All of this is achieved with a fairly narrow training corpus consisting mainly of $3d$ synthetic data from chaotic systems, which suggests there may be considerable room for further improvement by extending DynaMix’ training to include, e.g., various empirical time series.
One important take-home from this work therefore is that foundation models built based on principles of DS theory may be able to profoundly improve performance of current zero-shot TS forecasters. Most (if not all) empirically observed time series come from some underlying DS, and acknowledging this fact in model training and construction may help to advance the field. Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role. Most complex DS in nature Buzsaki (2006); Tziperman et al. (1997); Mumby et al. (2007), engineering Strogatz (2024), and society Mandelbrot and Hudson (2007); Watts and Strogatz (1998) are likely chaotic. Moreover, many chaotic attractors, like the Lorenz-63, feature a skeleton of infinitely many (unstable) periodic orbits of all possible periods Guckenheimer and Holmes (1983), thus inherently expressing a wide spectrum of temporal patterns.
@@
For now, we have mainly focused on rather stationary time series (but see Fig. 8b). Changes in statistical properties over time, tipping points, or widely differing time scales in the data, impose severe difficulties for zero-shot forecasting, as illustrated in Fig. 31a,b, which also plague other TS foundation models. This may partly be amended by explicitly including such types of non-stationary and multiscale DS in the training corpus. Another potential solution may be adding explicit filtering and decomposition modules, similar as in Auto- Wu et al. (2021) or FEDformers Zhou et al. (2022). A proof of concept of this idea is provided in Appx. A.5. Incorrect embeddings, which may not be immediately obvious, can also lead to forecast failures (Fig. 31c). Finally, continuous-time versions of the experts which can deal with sampling at irregular time points would be another fruitful direction.
@@ A.1 Methodological details
For training our model, we used a variant of sparse teacher forcing (STF), a control-theoretic method for DSR, with the theoretical background and rationale developed in Mikhaeil et al. [2022]. STF and related training methods Brenner et al. [2022, 2024b], Hess et al. [2023], Mikhaeil et al. [2022], Eisenmann et al. [2024] circumvent the exploding gradient problem when training on chaotic systems, while still enabling the DSR model to sufficiently forward-iterate trajectories into the future to capture long-term properties (see Mikhaeil et al. [2022] for details), and produces state-of-the-art results for DSR Brenner et al. [2024a, 2025]. In STF, current latent states states $\bm{z}_{t}$ are replaced with states inferred from the data at fixed intervals $\tau$ by (pseudo-)inversion of the decoder model, $\tilde{\bm{z}}_{t}=g^{-1}(\bm{x}_{t})$, thus recalibrating the trajectory at times ideally chosen based on the system’s Lyapunov spectrum Mikhaeil et al. [2022] (or simply treating $\tau$ as a hyper-parameter). In our case of an identity mapping from a latent subspace to the observations, $\hat{\bm{x}}_{t}=\mathcal{I}_{(N\times M)}\bm{z}_{t}$, this inversion becomes trivial, where $N$ is the number of observations (and thus readout neurons) and $\mathcal{I}_{(N\times M)}$ an $N\times M$ matrix with $\mathcal{I}_{(N\times M),rr}=1$ for $r\leq N$ and zeros else. Thus, during training, we force
@@
where we chose $\lambda=0.1$ and $c=0.01$. Rectified adaptive moment estimation (RADAM) Liu et al. [2020] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total. We used a learning rate exponentially decaying from $\eta_{\text{start}}=5\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$. The context length used in training was set to $T_{C}=500$, and the window of overlap with the model-generated time series to $\Delta t=50$, see Sect. 3.2. Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $30$ seconds depending on sequence length and model size. At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step. Hyperparameters for training were partly selected according to previous results in Brenner et al. [2024a, 2022], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.1,0.2\}$, $\Delta t=\{0,50,80\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
@@
We use $J=10$ AL-RNN experts for our model. Each expert has a latent dimension of $M=30$, of which $P=2$ are rectified-linear units (ReLUs). As suggested in Brenner et al. [2024a], we only use the first $N$ of the MPM-P linear units for the readout. We followed the initialization protocol in Brenner et al. [2022], Talathi and Vartak [2016], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix. The initial latent state was estimated as
@@
The gating network is implemented using a single-layer CNN with three channels and a kernel size of 2, stride of 1, and zero padding, with the identity as activation function. The MLP, which takes as inputs the concatenated weighted CNN outputs and current latent state, consists of two layers with ReLU activation. We initialize the temperature weights $\tau_{\text{att}}$ and $\tau_{\text{exp}}$ to $0.1$, the covariance matrix $\bm{\Sigma}=0.05\cdot\mathds{1}$, matrix $\bm{D}=\mathcal{I}_{(N\times M)}$, and draw all other CNN and MLP matrices from a Gaussian with standard deviation $0.01$. Model hyperparameters were selected according to previous results in Brenner et al. [2024a] and by probing a few relevant parameter settings, namely $J=\{1,3,5,10,20,30\}$ (Fig. 25), $\text{CNN-layers}=\{1,2\}$, $\text{CNN-kernel-size}=\{2,3,5\}$, $\text{MLP-layers}=\{1,2,3\}$, and different activation functions $\{\mathds{1},\tanh{},\text{ReLU}\}$ for the MLP/CNN. We found that an extensive grid search was not necessary to obtain a well-performing model.
@@
As in Koppe et al. [2019], Mikhaeil et al. [2022], Hess et al. [2023], Brenner et al. [2024b], Pals et al. [2024], Gilpin [2023], Zhang and Gilpin [2025], we assess the geometric (dis)similarity between true and model-generated attractors using a Kullback-Leibler divergence ($D_{\text{stsp}}$) defined across state space. Specifically, $D_{\text{stsp}}$ determines the overlap between the distributions of true trajectory points, $p_{true}(\bm{x})$, and of model-generated trajectories, $p_{gen}(\bm{x}|\bm{z})$, given by
@@
While especially for higher-dimensional state spaces these distributions may be approximated by Gaussian mixture models Koppe et al. [2019], here a discretized version was sufficient, with the state space parcellated into $K=m^{N}$ bins, where $m$ is the number of bins per dimension and $N$ the dimensionality of the system. We estimate the occupation probabilities $p_{i}$ of each bin via the relative frequencies $\hat{p}_{i}$ of trajectory visits and approximate $D_{\text{stsp}}$ as
@@
We set $m=30$ bins per dimension, following Hemmer et al. [2024]. For the $6d$ Lorenz-96 we use $m=5$ and for the empirical time series $m=20$. To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled from the DS.
To evaluate the agreement in long-term temporal dynamics, we compute the Hellinger distance $D_{H}$ between the power spectra of the true and model-generated time series Mikhaeil et al. [2022], Hess et al. [2023], defined as
@@
where $F(\omega)$ and $G(\omega)$ are the power spectra of the true and generated time series, respectively. Power spectra are obtained via the dimension-wise Fast Fourier Transform (FFT), smoothed using a Gaussian kernel (with $\sigma=20$ for DS and $\sigma=2$ for the short empirical time series), and normalized to ensure comparability. High-frequency tails dominated by noise are truncated as described in Hess et al. [2023]. The aggregated Hellinger distance $D_{H}$ is then computed as the average across all dimension-wise spectral comparisons.
Note that both $D_{\text{stsp}}$ and $D_{H}$ are assumed to assess (dis-)agreement in long-term properties in the limit $T\rightarrow\infty$. Hence, they are only sensible on rather long-term horizons, with $T=10,000$ used here.
For assessing short-term forecast quality aggregated across the whole test set of $54$ DS, we used a normalized $n$-step ahead prediction error (with $n=10$ in our evaluation) as recommended in Hewamalage et al. [2023], given by
@@
For individual comparisons on single empirical times, the MAE for evaluating short term predictions Hewamalage et al. [2023] was used:
@@
For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm Rosenstein et al. [1993]. First, the time series is embedded into a state space using time-delay embedding as in eq. 5 (e.g. using the PECUZAL algorithm Kraemer et al. [2021]). For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity Rosenstein et al. [1993]) and $l_{s}=0.25$. The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
@@
In chaotic systems, for a suitable choice of $k_{max}$ (here: $50$), this quantity initially grows approximately linearly with time Kantz and Schreiber [2004], with the slope given by $\lambda_{\max}>0$. In contrast, periodic (cyclic) systems have $\lambda_{\max}=0$. As for $D_{\text{stsp}}$ and $D_{H}$, Lyapunov exponents can only be reasonably assessed for sufficiently long time series, with practical guidelines of $T\approx 10^{d}-30^{d}$, where $d$ is the attractor’s fractal dimension Wolf et al. [1985], Rosenstein et al. [1993].